OpenStudy (anonymous):

How does this: (x^2-2x)+(y^2+4y)=7 become this: (x^2-2x+1) + (y^2 +4y+4)= 7+1+4?

OpenStudy (anonymous):

complete the square for x and y

OpenStudy (experimentx):

how does (x^2-2x+1) + (y^2 +4y+4)= 7+1+4 become (x^2-2x)+(y^2+4y)=7 ??

OpenStudy (anonymous):